An equipment has been purchased for a project to be completed in 6 years. The initial cost of equipment (6)
was Rs. 7,500,000. It is estimated that the used equipment can be sold in auction at a price of Rs. 1,200,000
after completion of the project. The expected/planned yearly usage of the equipment is as follows:
Year Equipment Usage (Hours)
1 140
2 170
3 230
4 80
5 110
6 70

Develop the yearly depreciation schedule of the equipment according to:
a) Sum-of-Years-Digit Method
b) Units-of-production Method
c) Double Declining Balance Method
Answer #1
Sum of years digit method Calculation Units of production methods Calculation Double declining balance method Calculation
Cost 7500000 =depreciation base*fraction of year 7500000 =deprecition base*(hours worked during current year/total hours estimated)    75,00,000.00 =WDV*double dimishing rate
Depreciation for year 1 1800000 =6300000*6/21 1102500 =6300000*(140/800)    24,99,750.00 =7500000*0.3334
Balance end of year 5700000 6397500    50,00,250.00
Depreciation for year 2 1500000 =6300000*5/21 1338750 =6300000*(170/800)    16,66,583.33 =5000250*0.3334
Balance end of year 4200000 5058750    33,33,666.68
Depreciation for year 3 1200000 =6300000*4/21 1811250 =6300000*(230/800)    11,11,111.10 =3333666.68*0.3334
Balance end of year 3000000 3247500    22,22,555.57
Depreciation for year 4 900000 =6300000*3/21 630000 =6300000*(80/800)      7,40,777.77 =2222555.57*0.3334
Balance end of year 2100000 2617500    14,81,777.80
Depreciation for year 5 600000 =6300000*2/21 866250 =6300000*(110/800)      4,93,876.54 =740777.77*0.3334
Balance end of year 1500000 1751250      9,87,901.26
Depreciation for year 6 300000 =6300000*1/21 551250 =6300000*(70/800)      3,29,267.49 =987901.26*0.3334
Balance end of year 1200000 1200000      6,58,633.77
Original cost 7500000
Salvage value 1200000
Usefullife 6 years
Straight line depreciation rate =100/6
16.66666667
Double dimishing rate =2*stratighline rate
=2*16.67
33.34
Sum of years digit method
Depreication base =original cost - salvage value
=7500000-1200000
6300000
Sum of years =1+2+3+4+5+6
21
Year 1 fraction rate =6/21
Yeas 2 =5/21
Year 3 =4/21
Year 4 =3/21
Year 5 =2/21
Year 6 =1/21
Units of production
Total estimated hours of production =140+170+230+80+110+70
800